# Robustly overfitting latents for flexible neural image compression

### v1

@@ 4 Experiments
Following (Yang et al., 2020), we run all experiments with temperature schedule $\tau(t)=\min(\exp\{-ct\},\tau_{max})$, where $c$ is the temperature rate determining how fast temperature $\tau$ is decreasing over time, $t$ is the number of train steps for the refinement of the latents and $\tau_{max}\in(0,1)$ determines how soft the latents start the refining procedure. Additionally, we refine the latents for $t=2000$ train iterations, unless specified otherwise. For more hyperparameter settings, see Section 4.6. Further, the experimental results are obtained from a pre-trained model trained with $\lambda=0.01$.
4.1 Implementation details
The pre-trained mean-scale hyperpriors are trained from scratch on the full-size CLIC 2020 Mobile dataset (Toderici et al., 2020), mixed with the ImageNet 2012 dataset (Russakovsky et al., 2015) with randomly cropped image patches taken of size $256\times 256$. For ImageNet, only images with a size larger than $256$ for height and width are used to prevent bilinear up-sampling that negatively affects the model performance. We use the architecture of (Minnen et al., 2018), except for the autoregressive part as a context model. Instead, we use the regular convolutional architecture of (Ballé et al., 2018). The model package for the mean-scale hyperprior is from CompressAI (Bégaint et al., 2020). During training, each model is evaluated on the Kodak dataset (Kodak, ). We run all models and methods on a single NVIDIA A100 GPU.
The models are trained with $\lambda=\{0.001,0.0025,0.005,$ 0.01,0.02,0.04,0.08}0.01,0.02,0.04,0.08\}0.01 , 0.02 , 0.04 , 0.08 }, with a batch size of 32 and Adam optimizer with a learning rate set to $1e^{-4}$. The models are trained for 2M steps, except for model $\lambda=0.001$, which is trained for 1M steps and model $\lambda=0.08$, which is trained for 3M steps. Further, the models for $\lambda=\{0.04,0.08\}$ are trained with $256$ hidden channels and the model for $\lambda=0.001$ is trained with $128$ hidden channels. The remaining models are trained with hidden channels set to $192$.
4.2 Baseline methods
@@
Figure 2: Performance plots of (a) True R-D Loss (b) Difference in loss (c) PSNR (d) BPP
4.3 Experimental results
Overall performance
Figure 2 shows the overall performance for the refinement of the latents on the Kodak dataset with method: STE, uniform noise, $\operatorname{atanh}$ and SSL. The true R-D loss in Figure 1(a) shows that STE has trouble converging and uniform noise quickly converges compared to $\operatorname{atanh}$ and SSL. We find that SSL outperforms all other methods, including $\operatorname{atanh}$ in terms of the lowest true R-D loss. Looking at the difference between the method loss and true loss Figure 1(b), we find that SSL quickly and smoothly finds a balance between the method and true loss, while $\operatorname{atanh}$ has a big difference in the beginning and more peaks before eventually converging. Additionally, uniform noise shows a big difference between the method loss and true loss, indicating that adding uniform noise overestimates its method loss compared to the true loss. For R-D curves, we refer to Appendix A.2.
Table 1: True R-D loss for different $\tau_{max}$ settings of: $\operatorname{atanh}(v)$, linear, cosine and SSL with $a=\frac{4}{3}$. Lowest R-D loss per column is marked with: $\downarrow$. Note that the function containing $\operatorname{atanh}$ is unnormalized.
Function \$\tau_{max}$ 0.2 0.4 0.6 0.8 1.0
$\cos^{2}(\frac{v\pi}{2})$ $0.7496$ $0.7414$ $0.7393$ $0.7387$ $0.7384$
$\exp{\operatorname{atanh}(v)}$ 00 $0.0002$ $0.0022$ $0.0033$ $0.0038$
$1-v$ (linear) $0130.0$ 00 $0.0003$ $0.0006$
Temperature sensitivity
Now that we have assessed the overall performance, we analyze the stability of each method. Table 1 represents the stability of $\operatorname{atanh}$ and the SGA+ methods, expressed in true R-D loss, for different $\tau_{max}$ settings for the temperature schedule. As can be seen, the most optimal setting is with $\tau_{max}=1$ for each of the SGA+ methods. $\operatorname{atanh}$ obtains equal loss for $\tau_{max}\in[0.4,0.5]$. In general, we find that the linear method of SGA+ is least sensitive to changes in $\tau_{max}$ and has equal loss between $\tau_{max}\in[0.7,1]$. To further examine the stability of the linear function compared to $\operatorname{atanh}$, we subtract the best $\tau_{max}$, column-wise, from the linear and $\operatorname{atanh}$ of that column. We now see that the linear function is not only least sensitive to changes in $\tau_{max}$, but overall varies little compared to the best $\tau_{max}$ settings of the other methods. While the SSL has the largest drop in performance when reducing $\tau_{max}$, it achieves the highest performance overall for higher values of $\tau_{max}$.

### v3

@@ 4 Experiments
Following Yang et al. (2020), we run all experiments with temperature schedule $\tau(t)=\min(\exp\{-ct\},\tau_{max})$, where $c$ is the temperature rate determining how fast temperature $\tau$ is decreasing over time, $t$ is the number of train steps for the refinement of the latents and $\tau_{max}\in(0,1)$ determines how soft the latents start the refining procedure. Additionally, we refine the latents for $t=2000$ train iterations, unless specified otherwise. See Section 4.2 for the hyperparameter settings.
Implementation details
We use two pre-trained hyperprior models to test SGA+, for both models we use the package from CompressAI Bégaint et al. (2020). The first model is similar to the one trained in Yang et al. (2020). Note, that the refining procedure needs the same storage and memory as theirs. In Appendix C, implementation details and results of this model can be found. All experiments in this section are run with a more recent hyperprior-based model which is based on the architecture of Cheng et al. (2020). Additionally, this model reduces the R-D loss for $\operatorname{atanh}$ on average by $7.6\%$ and for SSL by $8.6\%$, compared to the other model. The model weights can be retrieved from CompressAI Bégaint et al. (2020). The models were trained with $\lambda=\{0.0016,0.0032,0.0075,0.015,0.03,0.045\}$. The channel size is set to $N=128$ for the models with $\lambda=\{0.0016,0.0032,0.0075\}$, refinement of the latents on Kodak with these models take approximately 21 minutes. For the remaining $\lambda$’s channel size is set to $N=192$ and the refining procedure takes approximately 35 minutes. We perform our experiments on a single NVIDIA A100 GPU.
Baseline methods
@@
4.1 Overall performance
Figure 3(a) shows the rate-distortion curve, using image quality metric PSNR versus BPP, of the base model and for refinement of the latents on the Kodak dataset with method: STE, uniform noise, $\operatorname{atanh}$ and SSL. We clearly see how SSL outperforms all other methods. In Appendix 1(a), the results after $t=500$ iterations are shown, where the performance is more pronounced.
To show how each of the methods behave, we take a closer look at the performance plots shown in Figure 2. The refinement results are obtained from the pre-trained model, trained with $\lambda=0.0075$. The true R-D loss in Figure 2(a) shows that STE performs worse and has trouble converging, which is reflected in the R-D curve. Uniform noise quickly converges compared to $\operatorname{atanh}$ and SSL. We find that SSL outperforms all other methods, including $\operatorname{atanh}$ in terms of the lowest true R-D loss at all steps. Looking at the difference between the method loss and true loss Figure 2(b), we find that both SSL and $\operatorname{atanh}$ converge to 00. Yet, the initial loss difference is smaller and smoother for SSL, compared to $\operatorname{atanh}$. Additionally, uniform noise shows a big difference between the method and true loss, indicating that adding uniform noise overestimates its method loss compared to the true loss.
Tecnick and CLIC
To test how our method performs on other datasets, we use the Tecnick Asuni and Giachetti (2014) and CLIC dataset. We run baselines $\operatorname{atanh}$ and the base model and compare against SSL with $a=2.3$. Figure 3(b) shows the corresponding R-D curves on the Tecnick dataset. Note that Appendix B.3 contains the results of the CLIC dataset which shows similar behavior as on Tecnick. We find that both refining methods improve the compression performance in terms of the R-D trade-off. Additionally, our proposed method outperforms $\operatorname{atanh}$ and shows how it improves performance significantly at all points. The R-D plot after $t=500$ iterations for both datasets can be found in Appendix B.1. Note, that these results show even more pronounced performance difference.
BD-Rate and BD-PSNR
To evaluate the overall improvements of our method, we computed the BD-Rate and BD-PSNR Bjontegaard for Kodak, Tecnick and CLIC in Table 1. We observe that across the board SSL achieves an improvement in both BD-PSNR and BD-Rate. After $500$ steps, SSL achieves almost double the BD-Rate reduction as atanh. The difference between the two becomes smaller at $2000$ steps. This underlines the faster convergence behavior of the SSL method.
(a) Kodak
(b) Tecnick
(c) Semi-multi-rate
Figure 3: R-D performance for SSL on (a) Kodak with the baselines, (b) Tecnick with the base model and $\operatorname{atanh}$ and (c) Kodak for semi-multi-rate behavior with $\operatorname{atanh}$. Best viewed electronically.
Table 1: Pairwise Comparison between atanh and SSL of BD-PSNR and BD-Rate.
BD-PSNR (dB) BD-Rate (%)
500 steps 2000 steps 500 steps 2000 steps
Kodak Tecnick CLIC Kodak Tecnick CLIC Kodak Tecnick CLIC Kodak Tecnick CLIC
Base vs SSL 0.50 0.57 0.56 0.82 0.95 0.89 -10.30 -11.60 -13.18 -16.23 -18.77 -20.11
Base vs Atanh 0.26 0.28 0.31 0.69 0.79 0.78 -5.52 -5.91 -7.37 -13.82 -15.93 -17.70
Atanh vs SSL 0.24 0.28 0.26 0.14 0.16 0.11 -5.04 -5.97 -6.20 -2.86 -3.34 -2.76
4.2 Qualitative results
In Figure 4, we demonstrate the visual effect of our approach. We compare the original image with the compressed image from base model with $\lambda=0.0016$, refinement method $\operatorname{atanh}$, and SSL. For the image compressed by SSL, we observe that there are less artifacts visible in the overall image. For instance, looking at the window we see more texture compared to the base model and $\operatorname{atanh}$ method.
(a) Original Image
(b) Base Model
(0.19 BPP vs 25.75 PSNR)
(c) $\operatorname{atanh}$
(0.19 BPP vs 25.91 PSNR)
(d) SSL
(0.19 BPP vs 25.98 PSNR)
Figure 4: Qualitative comparison of a Kodak image from pre-trained model trained with $\lambda=0.0016$. Best viewed electronically.
Hyperparameter settings
Refinement of the latents with pre-trained models, similar to the one trained in Yang et al. (2020), use the same optimal learning rate of $0.005$ for each method. Refinement of the latents with the models of Cheng et al. (2020) use a $10$ times lower learning rate of $0.0005$. Following Yang et al. (2020), we use the settings for $\operatorname{atanh}$ with temperature rate $\tau_{max}=0.5$, and for STE we use the smaller learning rate of $0.0001$, yet STE still has trouble converging. Note that, we tuned STE just as the other baselines. However, the STE method is the only method that has a lot of trouble converging. Even with smaller learning rates, the method performed poorly. The instability of training is not only observed by us, but is also observed in Yang et al. (2020); Yin et al. (2019). For SGA+, we use optimal convergence settings, which are a fixed learning rate of $0.0005$, and $\tau_{max}=1$. Experimentally, we find approximately best performance for SLL with $a=2.3$.
5 Analysis and Societal Impact
In this section we perform additional experiments to get a better understanding of SGA+. An in-depth analysis shows the stability of each proposed method, followed by an experiment that expresses changes in the true R-D loss performance when one interpolates between functions. Further, we evaluate three-class rounding for each of our methods. Finally, we show how SGA+ for semi-multi-rate behavior improves the performance of its predecessor and we discuss the societal impact. Note, the results of the additional experiments are obtained from the model trained with $\lambda=0.0075$.
Table 2: True R-D loss for different $\tau_{max}$ settings of: $\operatorname{atanh}(v)$, linear, cosine and SSL with $a=2.3$. The lowest R-D loss per column is marked with: $\downarrow$. Note that the function containing $\operatorname{atanh}$ is unnormalized.
Function \$\tau_{max}$ 0.2 0.4 0.6 0.8 1.0
$\exp{\operatorname{atanh}(v)}$ $0.6301$ $0.6273$ $0.6267$ $0.6260$ $0.6259$
$1-v$ (linear) $0.6291\downarrow$ $0.6229\downarrow$ $0.6225$ $0.6222$ $0.6220$
$\cos^{2}(\frac{v\pi}{2})$ $0.6307$ $0.6233$ $0.6194\downarrow$ $0.6186$ $0.6187$
$\sigma(-a\sigma^{-1}(v))$ $0.6341$ $0.6233$ $0.6196$ $0.6181\downarrow$ $\textbf{0.6175}\downarrow$
$\exp{\operatorname{atanh}(v)}$ $0.0010$ $0.0044$ $0.0073$ $0.0079$ $0.0084$
$1-v$ (linear) 00 00 $0.0031$ $0.0041$ $0.0045$
Temperature sensitivity
Table 2 represents the stability of $\operatorname{atanh}$ and the SGA+ methods, expressed in true R-D loss, for different $\tau_{max}$ settings for the temperature schedule. As can be seen, the most optimal setting is around $\tau_{max}=1$ for each of the methods. In the table we find that the linear function is least sensitive to changes in $\tau_{max}$. To further examine the stability of the linear function compared to $\operatorname{atanh}$, we subtract the best $\tau_{max}$, column-wise, from the linear and $\operatorname{atanh}$ of that column. Also taking into account the sensitivity results of the one of Yang et al. (2020) in Appendix C.2 we find in general, that overall performance varies little compared to the best $\tau_{max}$ settings of the other methods and has reasonable performance. While SSL has the largest drop in performance when reducing $\tau_{max}$, it achieves the highest performance overall for higher values of $\tau_{max}$. If there is no budget to tune the hyperparameter of SGA+, the linear version is the most robust choice. Further, we evaluated the necessity of tuning both the latents and hyper-latents. When only optimizing the latents with the linear approach for $\tau_{max}=1$, we found a loss of $0.6234$. This is a difference of $0.0012$, which implies that optimizing the hyper-latent aids the final loss.
Table 3: True R-D loss of two- versus three-class rounding for SGA+ with the extended version of the linear, cosine, and SSL method at iteration 2000 and in brackets after 500 iterations.
[…]
$f_{3c,\mathrm{linear}}(w|r=0.98,n=2.5)$ $0.6220$ ($0.6594$) $0.6175$($0.6435$)
$f_{3c,\mathrm{cosine}}(w|r=0.98,n=3)$ $0.6187$ ($0.6516$) $0.6175$ ($0.6449$)
$f_{3c,\mathrm{sigmoidlogit}}(w|r=0.93,n=2.5)$ $0.6175$ (0.6445) $0.6203$ (0.6360)
Interpolation
Table 4 represents the interpolation between different functions, expressed in true R-D loss. In Appendix B.2 the corresponding loss plots can be found. Values for $a<1$ indicate methods that tend to have larger gradients for $v$ close to the corners, while high values of $a$ represent a method that tends to a (reversed) step function. The smallest $a=0.01$ diverges and results in a large loss value compared to the rest. For $a=2.3$ we find a lowest loss of $0.6175$ and for $a=5$ we find fastest convergence compared to the rest. Comparing these model results with the model results with the one of Yang et al. (2020), see Appendix C.2, we find that the Cheng et al. (2020) model obtains more stable curves.
a R-D Loss
0.01 $0.7323$
0.3 $0.6352$
$0.65$ (approx atanh) $0.6260$
0.8 $0.6241$
$1$ (linear) $0.6220$
1.33 $0.6199$
$1.6$ (approx cosine) $0.6186$
2.3 $0.6175\downarrow$
5 $0.6209$
Three-class rounding
In Table 3, the true R-D loss for two versus three-class rounding can be found at iteration $t=2000$ and in brackets $t=500$ iterations. For each method, we performed a grid search over the hyperparameters $r$ and $n$. As can be seen in the table, most impact is made with the extended version of the linear of SGA+, in terms of the difference between the two versus three-class rounding at iteration $t=2000$ with loss difference $0.0045$ and $t=500$ with $0.0159$ difference. In Appendix B.3 a loss plot of the two- versus three- class rounding for the extended linear method can be found. Concluding, the three-class converges faster. For the extended cosine version, there is a smaller difference and for SSL we find that the three-class extension only boosts performance when run for $t=500$. In Appendix C.2, the three-class experiments for the pre-trained model similar to Yang et al. (2020) can be found. We find similar behavior as for the model of Cheng et al. (2020), whereas with the linear version most impact is made, followed by the cosine and lastly SSL. The phenomenon that three-class rounding only improves performance for $t=500$ iterations for SSL may be due to the fact that SSL is already close to optimal. Additionally, we ran an extra experiment to asses what percentage of the latents are assigned to the 3-classes. This is run with the best settings for the linear version $f_{3c,\mathrm{linear}}(w|r=0.98,n=2.5)$. At the first iteration, the probability is distributed as follows: $p(y=\lfloor v\rceil)=0.9329$, for $p(y=\lfloor v\rceil-1)=0.0312$, and $p(y=\lfloor v\rceil+1)=0.0359$. This indicates that the class probabilities are approximately $3.12$ for class $-1$ and $3.6$ for class $+1$. This is a lot when taking into account that many samples are taken for a large dimensional latent. In conclusion, three-class rounding may be attractive under a constraint optimization budget, possibly because it is easier to jump between classes. Additionally, for the extended linear three-class rounding also results in faster convergence.
Semi-multi-rate behavior
An interesting observation is that one does not need to use the same $\lambda$ during refinement of the latents, as used during training. This is also mentioned in Gao et al. (2022) for image and Xu et al. (2023) for video compression. As a consequence of this approach, we can optimize to a neighborhood of the R-D curve without the need to train a new model from scratch. We experimented and analyze the results with methods: $\operatorname{atanh}$ and SSL. Figure 3(c) shows the performance after $t=500$ iterations for the model of Cheng et al. (2020). We find that SSL is moving further along the R-D curve compared to $\operatorname{atanh}$. Note the refinement does not span the entire curve and that the performance comes closer together for running the methods longer, see Appendix B.4. For future work it would be interesting how SGA+ compares to the methods mentioned in Gao et al. (2022); Xu et al. (2023), since SSL outperforms $\operatorname{atanh}$.
Societal impact
The improvement of neural compression techniques is important in our data-driven society, as it allows quicker development of better codecs. Better codecs reduce storage and computing needs, thus lowering costs. However, training these codes requires significant computational resources, which harms the environment through power consumption and the need for raw materials.
6 Conclusion and Limitations
In this paper we proposed SGA+, a more effective extension for refinement of the latents, which aids the compression performance for pre-trained neural image compression models. We showed how SGA+ has improved properties over SGA and we introduced SSL that can approximately interpolate between all of the proposed methods. Further, we showed how our best-performing method SSL outperforms the baselines in terms of the R-D trade-off and how it also outperforms the baselines on the Tecnick and CLIC dataset. Exploration of SGA+ showed how it is more stable under varying conditions. Additionally, we gave a general notation and demonstrated how the extension to three-class rounding improves the convergence of the SGA+ methods. Lastly, we showed how SGA+ improves the semi-multi-rate behavior over SGA. In conclusion, especially when a limited computational budget is available, SGA+ offers the option to improve the compression performance without the need to re-train an entire network and can be used as a drop-in replacement for SGA.
Besides being effective, SGA+ also comes with some limitations. Firstly, we run each method for 2000 iterations per image. In practice this is extremely long and time consuming. We find that running the methods for 500 iterations already has more impact on the performance and we would recommend doing this, especially when a limited computational budget is available. Future work may focus on reducing the number of iterations and maintaining improved performance. Note, higher values for $a$ flatten out quickly, but they achieve much better gains with low-step budgets. Further, the best results are obtained while tuning the hyperparameter of SSL and for each of our tested models this lead to different settings. Note, that the experiments showed that the linear version of SGA+ is least sensitive to hyperparameter changes and we would recommend using this version when there is no room for tuning. Additionally, although three-class rounding improves the compression performance in general, it comes with the cost of fine-tuning extra hyperparameters. Finally, it applies for each method that as the temperature rate has reached a stable setting, the performance will be less pronounced, the longer you train, but in return requires extra computation time at inference.
Acknowledgments
This work was carried out on the Dutch national e-infrastructure with the support of SURF Cooperative.
Appendix A Proof
In this appendix we will proof that the normalization from the (Gumbel) softmax causes infinite gradients at 0.
Recall that the probability is given by a 2-class softmax is defined by:
$K(v)=\frac{e^{f(v)}}{e^{f(v)}+e^{g(v)}},$
where $f(v)=-\text{atanh}(v)$ and $g(v)=-\text{atanh}(1-v)$. We will study the softmax for the first class 00, since the softmax is symmetric this also holds for the second class. The problem is that the gradients of the function $K(v)$ will tend to $\infty$ for both $v\to 1$ but also for $v\to 0$. Here we show that the gradients also tend $\infty$ for $v\to 0$, via the normalization with the term $g(v)$. First, take the derivative to $v$:
$\frac{dK(v)}{dv}=\frac{dK(v)}{df(v)}\cdot\frac{df(v)}{dv}+\frac{dK(v)}{dg(v)}\cdot\frac{dg(v)}{dv},$
where $\frac{dK(v)}{df(v)}=K(v)(1-K(v))$ and $\frac{dK(v)}{dg(v)}=-K(v)\frac{e^{g(v)}}{e^{f(v)}+e^{g(v)}}$. Recall that $\frac{d\text{atanh}(v)}{dv}=\frac{1}{1-v^{2}}$ therefore $\frac{df(v)}{dv}=-\frac{1}{1-v^{2}}$ and $\frac{dg(v)}{dv}=\frac{1}{1-(1-v)^{2}}$. Plugging this in and computing $\frac{dK(v)}{dg(v)}$ gives us:
$\frac{dK(v)}{dv}=K(v)(1-K(v))\left(-\frac{1}{1-v^{2}}\right)-K(v)\cdot\frac{e^{g(v)}}{e^{f(v)}+e^{g(v)}}\cdot\frac{1}{1-(1-v)^{2}}.$
Taking the limit to 00, (recall that $\lim_{v\to 0}K(v)=1,\lim_{v\to 0}e^{f(v)}=1$ and $\lim_{v\to 0}e^{g(v)}=0$) allows the following simplifications:
$\lim_{v\to 0}0\cdot\left(-\frac{1}{1-0^{2}}\right)-1\cdot\frac{e^{g(v)}}{1+0}\cdot\frac{1}{1-(1-v)^{2}}$
For simplicity we substitute $q=1-v$ (when $q\to 1$, then $v\to 0$) which will result in the following:
$\lim_{v\to 0}-e^{-\text{atanh}(1-v)}\cdot\frac{1}{1-(1-v)^{2}}=\lim_{q\to 1}-e^{-\text{atanh}(q)}\cdot\frac{1}{1-q^{2}},$
Recall, $-\text{atanh}(q)=-\frac{1}{2}\ln{\frac{1+q}{1-q}}$, so $e^{-\text{atanh}(q)}=1/\sqrt{\frac{1+q}{1-q}}$ thus:
$-\lim_{q\to 1}\sqrt{\frac{1-q}{1+1}}\cdot\frac{1}{1-q^{2}}=-\lim_{q\to 1}\sqrt{\frac{1}{2}\frac{(1-q)}{(1-q^{2})^{2}}}$
Since $\frac{1}{2}$ is a constant and $\lim_{x\to\infty}\sqrt{x}=\infty$ the final step is to simplify and solve:
$-\lim_{q\to 1}\sqrt{\frac{(1-q)}{(1-q^{2})^{2}}}=-\lim_{q\to 1}\sqrt{\frac{-1}{(q-1)(q+1)^{2}}}=-\infty.$
This concludes the proof that the gradients tend to $\infty$- ∞ for $v\to 0$.
Appendix B Additional results
In this appendix, additional experimental results for refinement of the latents with the pre-trained models of Cheng et al. (2020) can be found.
Difference across runs
Although we do not report the standard deviation for every run, the difference across runs is very small. For the model of Cheng et al. (2020) with $\lambda=0.0032$, running five refinement procedures for 2000 iterations, results in a mean of $0.3969$ and a standard deviation of $2.41\cdot 10^{-5}$.
B.1 Additional overall performance
Figure B.1 show the R-D curve for the Kodak dataset. We observe that $\operatorname{atanh}$ is similar to uniform noise at $t=500$ iterations, while SSL manages to achieve better R-D gains. After $t=2000$ iterations SLL achieves the best R-D trade-off, but the gain is a bit smaller compared to $\operatorname{atanh}$ at $t=500$ iterations.
Tecnick and CLIC
Figure B.2 and Figure B.3 show the R-D results for respectively, Tecnick and CLIC at $t=\{500,2000\}$ iterations. We observe that SSL achieves best performance at $t=2000$, compared to $t=500$ iterations. However, running the method for $2000$ iterations is in practice very long. Looking at the results at $t=500$ iterations we see that there is already great improvement of performance for SSL, compared to the base model for both Tecnick and CLIC dataset.
(a) Kodak ($t=500$)
(b) Kodak ($t=2000$)
Figure B.1: Comparison of $\operatorname{atanh}$ and SSL on the Kodak dataset for $t=\{500,2000\}$ iterations.
(a) Tecnick ($t=500$)
(b) Tecnick ($t=2000$)
Figure B.2: Comparison of $\operatorname{atanh}$ and SSL on the Tecnick dataset for $t=\{500,2000\}$ iterations.
(a) CLIC ($t=500$)
(b) CLIC ($t=2000$)
Figure B.3: Comparison of $\operatorname{atanh}$ and SSL on the CLIC dataset for $t=\{500,2000\}$ iterations.
B.2 Interpolation
In Figure 4(a) the true loss, difference in loss, BPP and PSNR curves can be found. As can be seen for $a=0.01$, the function diverges and $a=0.3$ does not seem to reach stable behavior, both resulting in large loss values. For $a\geq 1$, the difference in losses start close to zero (see Figure 7(b)). SSL with $a=5$ results in the fastest convergence and quickly finds a stable point but ends at a higher loss than most methods.
@@ B.2 Interpolation
Figure B.4: Interpolation performance plots of different $a$ settings for SLL (a) True R-D Loss (b) Difference in loss (c) PSNR (d) BPP.
B.3 Two- versus three- class rounding
Besides an improved RD performance for the three-class rounding, we also found that three-class rounding leads to faster convergence. In Figure B.5 a true loss plot over the iterations for the linear method, can be found. As one can see, the three-class converges faster and especially on 500 iterations, boosts performance which makes it attractive under a constraint budget.
Figure B.5: True R-D loss curves for two- versus three-class rounding of the linear method.
B.4 Semi-multi-rate behavior
In Figure B.6, we have plotted the R-D curve of the base model (lime green line) and its corresponding R-D curves, obtained when refining the latents with the proposed $\lambda$’s. For each model trained using $\lambda\in\{0.0016,0.0032,0.0075,0.015,0.03,0.045\}$, we run $\operatorname{atanh}$ and SSL with $a=2.3$ for $t=2000$ iterations for all $\lambda\in\{0.0004,0.0008,0.0016,0.0032,0.0075,0.015,0.03,0.045,0.06,0.09\}$. We depicted the base curve alongside the curves for each base model. We observe that the improved performance by SSL is especially noticeable at $t=500$ iterations in Figure 6(a).
(a) $t=500$ iterations
(b) $t=200$ iterations
Figure B.6: R-D performance on Kodak of the Cheng et al. (2020) model when varying the target $\lambda$. Best viewed electronically.
Appendix C Mean-Scale Hyperprior
To make a clear comparison we trained a similar mean-scale hyperprior as in Yang et al. (2020). Therefore, we use the architecture of Minnen et al. (2018), except for the autoregressive part as a context model. Instead, we use the regular convolutional architecture of Ballé et al. (2018). The model package for the mean-scale hyperprior is from CompressAI Bégaint et al. (2020). The details and results of this model can be found in this section.
Similar as for Cheng et al. (2020) we run all experiments with temperature schedule $\tau(t)=\min(\exp\{-ct\},\tau_{max})$. Additionally, we refine the latents for $t=2000$ train iterations, unless specified otherwise.
Table C.1: True R-D loss results for the interpolation between different functions by changing $a$ of the SSL.